# A Training-Free Length Extrapolation Approach for LLMs: Greedy Attention Logit Interpolation (GALI)

### v1

@@ Abstract
Transformer-based Large Language Models (LLMs) struggle to process inputs exceeding their training context window, with performance degrading due to positional out-of-distribution (O.O.D.) that disrupt attention computations. Existing solutions, fine-tuning and training-free methods, are limited by computational inefficiency, attention logit outliers or loss of local positional information. To address this, we propose Greedy Attention Logit Interpolation (GALI), a training-free length extrapolation method that maximizes the utilization of pretrained positional intervals while avoiding attention logit outliers through attention logit interpolation. The result demonstrates that GALI consistently outperforms state-of-the-art training-free methods. Our findings reveal that LLMs interpret positional intervals unevenly within their training context window, suggesting that extrapolating within a smaller positional interval range yields superior results—even for short-context tasks. GALI represents a significant step toward resolving the positional O.O.D. challenge, enabling more reliable long-text understanding in LLMs. Our implementation of GALI, along with the experiments from our paper, is open-sourced at https://github.com/AcademyCityL/GALI.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) have become indispensable for a wide range of natural language processing tasks, yet their performance is fundamentally constrained by the training context window, i.e., the maximum input length used during training. When tasked with processing input text that exceeds this predefined limit, LLMs exhibit sharp performance degradation, with perplexity (PPL) increasing exponentially as input length grows (Peng et al., 2023; Chen et al., 2023b; Xiao et al., 2023; Han et al., 2024; Jin et al., 2024; An et al., 2024b). This limitation poses significant challenges for applications requiring robust long-text understanding, such as document summarization, legal text analysis, and conversational AI.
@@
The core issue lies in the model’s inability to generalize beyond the positional distributions encountered during pretraining, leading to disruptions in attention score computations—a phenomenon known as positional out-of-distribution (O.O.D.) (Chen et al., 2023b; Jin et al., 2024; Xu et al., 2024). Addressing positional O.O.D. is critical for enhancing LLMs’ length extrapolation capabilities and enabling reliable long-text processing.
@@
Existing approaches to mitigating positional O.O.D. can be classified into three categories: (1) Lambda-Shaped Attention Mechanisms, which stabilize PPL but compromise the ability to capture long-range dependencies across distant tokens (Xiao et al., 2023; Han et al., 2024; Jiang et al., 2024; Li et al., 2024a; Zhang et al., 2024); (2) Fine-Tuning on long texts, which involves training on datasets with extended positional contexts using interpolation (Chen et al., 2023b, a; Xiong et al., 2023; Li et al., 2023; Ding et al., 2024; Li et al., 2024b; Wu et al., 2024) or extrapolation (Zhu et al., 2023; Chen et al., 2023c; Ding et al., 2024). While effective, this approach is resource-intensive and still encounters cases where positional IDs exceed its fine-tuned context window; and (3) Training-free length extrapolation methods, which include Rotary Position Embedding (RoPE) frequency interpolation techniques (e.g., Neural Tangent Kernel (NTK), Dyn-NTK, YaRN) (LocalLLaMA, 2023b, a; Peng et al., 2023) and inputs rearrangement strategies (e.g., SelfExtend, ChunkLlama) (Jin et al., 2024; An et al., 2024b). They partially mitigate positional O.O.D. but suffer from limitations, such as the emergence of attention logit outliers or the loss of local positional information.
To overcome the challenges above, this paper introduces Greedy Attention Logit Interpolation (GALI), a novel training-free length extrapolation method tailored for transformer-based LLMs. GALI has two primary objectives: (1) to maximize the utilization of pretrained positional intervals, thereby minimizing disruption to attention computations, and (2) to mitigate outliers in attention logits through an attention logit interpolation strategy.
Key Innovations in GALI: First, GALI optimizes the utilization of pretrained position intervals. During the prefill phase, the method retains the model’s native position IDs for tokens within the training context window, ensuring that the corresponding hidden states maintain the quality achievable by the original model. GALI assigns interpolated position IDs chunk-wise for tokens exceeding the training context window, balancing efficiency with quality. During the decode phase, GALI dynamically generates interpolated position IDs for newly generated tokens, minimizing interpolation-induced disruptions during token generation. Second, GALI introduces a novel approach to compute attention logits for interpolated positional intervals. With the help of linear interpolated position IDs, GALI performs local linear interpolation directly on attention logits instead of computing positional embeddings, avoiding the outlier issues commonly associated with positional O.O.D. Furthermore, it simulates the oscillatory characteristics of RoPE, by adding Gaussian noise proportional to relative distances, enhancing the model’s positional sensitivity.
GALI offers a training-free, plug-and-play solution that operates exclusively during inference by maximizing the utilization of pretrained positional intervals and focusing on attention logit interpolation. It extends the usable context window of LLMs, significantly improving long-text processing capabilities while integrating seamlessly with existing long-context frameworks.
Empirical Evaluation: We evaluated GALI across three benchmarks, alongside a quantitative analysis of attention distribution, covering both real-world long-context applications and long-context language modeling. The results demonstrate that GALI consistently outperforms state-of-the-art training-free methods, achieving superior performance across diverse long-context scenarios. Furthermore, our analysis reveals a performance-enhancing strategy based on inconsistencies in how models interpret positional intervals. Notably, applying extrapolation methods within a smaller range of positional intervals leads to better results, even for short-context tasks. In summary, GALI represents a significant advancement in addressing the positional out-of-distribution (O.O.D.) challenge, unlocking new potential for LLMs in long-text understanding and processing.
Our main contributions are summarized as follows:
  • A Novel Training-Free Length Extrapolation Method: We introduce Greedy Attention Logit Interpolation (GALI), a training-free method that maximizes the utilization of pretrained positional intervals while leveraging attention logit interpolation to mitigate outliers, effectively addressing the positional O.O.D. problem in long-text processing for LLMs.
  • Comprehensive Evaluation and Analysis: We conducted a comprehensive comparison and analysis of GALI and other state-of-the-art training-free length extrapolation methods. While demonstrating GALI’s superior performance, we also identified a more effective length extrapolation strategy.
2 Background
@@
Rotary Position Embedding (RoPE): RoPE (Su et al., 2024) is a technique that encodes positional information by applying rotary transformations to token embeddings, enabling relative position modeling in transformers. Given two token embeddings $\bm{x}_{m},\bm{x}_{n}\in\mathbb{R}^{l}$ as query and key corresponding to position $m$ and $n$, the projection matrix $\bm{W}_{Q},\bm{W}_{K}\in\mathbb{R}^{d\times l}$, RoPE applies a rotation to the projected token embeddings, i.e., $\bm{q}_{m}^{r}=(\bm{W}_{Q}\bm{x}_{m})e^{im\bm{\theta}}$, $\bm{k}_{n}^{r}=(\bm{W}_{K}\bm{x}_{n})e^{in\bm{\theta}}$, where $\bm{\theta}=[b^{0},b^{-2/d},\dots,b^{-2(j-1)/d}]$, $j\in[1,2,\dots,d/2]$ and $b$ is originally set to 10000. After that, the inner product between the query $\bm{q}_{m}^{r}$ and key $\bm{k}_{n}^{r}$ can be represented by the real part of ${\bm{q}_{m}^{r}}^{*}\bm{k}_{n}^{r}$, i.e.:
@@
$g(\cdot)$ is the function mapping token embeddings $\bm{x}_{m},\bm{x}_{n}$ to the attention logit, which depends on their relative distance and is irrelevant to their absolute positions. Additionally, RoPE exhibits a long-term decay as relative distance increases (Su et al., 2024), as illustrated in Figure 1. Our proposed method, GALI, leverages two key properties of RoPE to achieve position interpolation and length extrapolation effectively.
Figure 1: Visualization of long-term decay in attention logits. The sentence “The quick brown fox jumps over the lazy dog.” is fed into a one-layer Llama3-8b-ins model. Figure 1 shows the attention logits from the last token to the second-to-last token. Figure 1 presents the logits from the last token to the first token. As the positional ID interval increases from 1 to 819, a clear decay phenomenon in the logits is observed.
Positional Out-Of-Distribution (O.O.D.): In Transformer architectures, the self-attention mechanism is inherently position-agnostic, necessitating the use of position embeddings to encode positional information for processing ordered inputs (Dufter et al., 2021; Kazemnejad et al., 2024). Even in large language models (LLMs) with causal attention, explicit positional encoding through position embeddings remains the standard approach. [^1] Recent studies suggest causal attention implicitly encodes positional information, enabling performance without explicit position embeddings. However, this is beyond the scope of this paper. During inference, when LLMs encounter input sequences exceeding the maximum length seen during training, the use of unseen position IDs causes a positional out-of-distribution (O.O.D.) issue, leading to degraded performance (Chen et al., 2023b; Jin et al., 2024; Xu et al., 2024). In the Rotary Position Embedding (RoPE) mechanism, extrapolating position IDs beyond the training range introduces untrained positional intervals, disrupting the attention score distribution (Chen et al., 2023b). In contrast, position interpolation has yielded more stable attention distributions, requiring fewer fine-tuning steps (Chen et al., 2023b). This observation has inspired subsequent interpolation-based methods (Chen et al., 2023b, a; Xiong et al., 2023; Li et al., 2023; Ding et al., 2024; Li et al., 2024b; Wu et al., 2024), as well as training-free approaches that map position interpolation into alternative frequency dimensions in embeddings (LocalLLaMA, 2023b, a; Peng et al., 2023). Recent work has explored other training-free length extrapolation techniques, such as group position IDs (Jin et al., 2024) or chunk attention (An et al., 2024b).
@@ 3 Method
Figure 2: The overall procedure of the proposed GALI framework. The green dashed line illustrates position ID interpolation, while the blue dashed line shows attention logit interpolation. In this example, the training context window $L_{tr}$ is 4, the chunk size $s$ is 2, the local window $L_{w}$ is 2, and the prefill length is 6. Chunks are denoted as $c_{1}$ (first chunk) and $c_{2}$ (second chunk), while $t_{7}$ represents the first generated token. The relative distance matrix $R$ incorporates $\lceil R\rceil$, $\lfloor R\rfloor$, and $R^{\prime}$, representing ceiling, floor, and modulo operations, respectively. Red numbers in the relative distance matrix indicate interpolated positional intervals and $\mathcal{N}$ represents the Gaussian noise.
This section presents Greedy Attention Logit Interpolation (GALI), a novel training-free position interpolation method designed to enhance the utilization of pretrained positional information while ensuring stable attention computations. GALI is guided by two objectives: (1) Maximizing the use of pretrained positional information, i.e., the positional intervals encountered during pre-training. To do so, we perform minimal-impact position interpolation for each chunk in the prefill phase and newly generated tokens during the decoding phase. (2) Preventing outliers in attention logits by avoiding direct computation of position embeddings for interpolated positions. We apply local linear interpolation to the attention logits based on interpolated position IDs and simulate RoPE’s oscillatory behavior by introducing Gaussian noise. The process is shown in Figure 2.
@@ 3.1 Position ID Interpolation
The proposed GALI introduces a strategy for position ID interpolation to handle inputs exceeding the maximum training length, aiming to fully utilize pretrained relative positional intervals. During the prefill phase, GALI segments the input portion beyond the maximum training length into multiple chunks and performs position ID interpolation for each chunk individually. In contrast, the portion within the maximum training length remains unchanged, without any interpolation. This approach provides several advantages. First, if the input length does not exceed the maximum training length, the prefill phase remains entirely unaffected, preserving the model’s original capabilities. Second, when the input length surpasses the maximum training length, the text within the training range can fully leverage the pretrained relative positional intervals to generate hidden states based on the model’s original performance. Finally, for inputs beyond the training length, GALI can apply position ID interpolation in a manner that minimizes the impact on model quality for each chunk. Specifically, each chunk only adds the exact number of new position IDs it requires. This ensures the interpolation process maintains a balance between computational efficiency and fidelity.
This approach draws inspiration from Dyn-NTK(LocalLLaMA, 2023a), which adjusts scaling factors to handle inputs of varying lengths. However, dynamic NTK applies a uniform interpolation across the entire input during the prefill phase, potentially overlooking the varying interpolation requirements of individual tokens. For instance, a token just beyond the training context window only requires the interpolation of one new position ID, whereas subsequent tokens would require progressively more. Ideally, the interpolation for each token would be customized to minimize its impact on quality. However, performing interpolation at the token level is computationally inefficient. To address these challenges, GALI segments the input exceeding the training length into chunks and applies interpolation only for the position IDs that extend beyond the training range within each chunk. By tailoring interpolation to the specific needs of each chunk, GALI avoids the uniform scaling factors used in previous methods, such as NTK, Dyn-NTK, YaRN, or SE, which can lead to suboptimal results. This chunk-wise strategy ensures a more adaptive and efficient interpolation process while maintaining model performance.
@@
Building on insights from prior studies, GALI retains a local window of size $L_{w}$ around the current token to preserve the original positional intervals within this range. This design allows the current token to effectively interpret its immediate context using the model’s pretrained capabilities. During the decode phase, where tokens are generated sequentially, GALI dynamically interpolates the required position IDs for each newly generated token, focusing on minimizing quality degradation. Specifically, in the prefill stage, given an input sequence $S=(w_{1},w_{2},\dots,w_{L_{tr}},\dots,w_{L})$ where $L_{tr}$ is the size of training context window and $L$ is the input length in the prefill stage. We first divide it into chunks $C=(c_{1},c_{2},\dots,c_{L_{c}})$, where the size of $c_{1}$ is $L_{tr}$ and other chunks have a size $s$, so the $L_{c}=\lceil{\frac{L-L_{tr}}{s}}\rceil$. After that, we assign position IDs $S=(s_{1},s_{2},\dots,s_{L_{c}})$ for each chunk, where $s_{1}=[0,1,\dots,L_{tr}-1]$ and others are interpolated position IDs according to the following formula:
$\begin{cases}&j>1;\\ &{g}_{j}=\lceil{\frac{\sum_{i=1}^{j}{len(c_{i})}-L_{w}}{L_{tr}-L_{w}}}\rceil;\\ &{v}_{j}=1/{g}_{j};\\ &s_{j}=[0,1*{v}_{j},2*{v}_{j},\dots,({g}_{j}-1)*{v}_{j},1,\dots,\\ &L_{tr}-L_{w}-1,L_{tr}-L_{w},\dots,L_{tr}-1]\end{cases}$ (2)
We first calculate the minimal group size ${g}_{j}$ required for each chunk after excluding the local window $L_{w}$ and then apply linear interpolation within each group. During the decode phase, each newly generated token is treated as an individual chunk of size 1, enabling the computation of minimally impactful interpolated position IDs.

### v2

@@ Abstract
Transformer-based Large Language Models (LLMs) struggle with inputs exceeding their training context window due to positional out-of-distribution (O.O.D.) issues that disrupt attention. Existing solutions, including fine-tuning and training-free methods, face challenges like inefficiency, redundant interpolation, logit outliers, or loss of local positional information. We propose Greedy Attention Logit Interpolation (GALI), a training-free method that improves length extrapolation by greedily reusing pretrained positional intervals and interpolating attention logit to eliminate outliers. GALI achieves stable and superior performance across a wide range of long-context tasks without requiring input-length-specific tuning. Our analysis further reveals that LLMs interpret positional intervals unevenly and that restricting interpolation to narrower ranges improves performance, even on short-context tasks. GALI represents a step toward more robust and generalizable long-text processing in LLMs. Our implementation of GALI, along with the experiments from our paper, is open-sourced at https://github.com/adlnlp/Gali.
A Training-Free Length Extrapolation Approach for LLMs:
Greedy Attention Logit Interpolation
Yan Li1, Tianyi Zhang2, Zechuan Li3, Soyeon Caren Han2thanks: Corresponding Author 1The University of Sydney, 2The University of Melbourne, 3Hunan University 1yali3816@uni.sydney.edu.au, caren.han@unimelb.edu.au
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) have become indispensable for a wide range of natural language processing tasks, yet their performance is fundamentally constrained by the training context window, i.e., the maximum input length used during training. When tasked with processing input text that exceeds this predefined limit, LLMs exhibit sharp performance degradation, with perplexity (PPL) increasing exponentially as input length grows Xiao et al. (2023); Han et al. (2024). This limitation poses significant challenges for applications requiring robust long-text understanding, such as document summarization, legal text analysis, and conversational AI.
@@
The core issue lies in the model’s inability to generalize beyond the positional distributions encountered during pretraining, leading to disruptions in attention score computations, a phenomenon known as positional out-of-distribution (O.O.D.) Chen et al. (2023b); Jin et al. (2024); Xu et al. (2024). Addressing positional O.O.D. is critical for enhancing LLMs’ length extrapolation capabilities and enabling reliable long-text processing.
@@
Existing approaches to mitigating positional O.O.D. can be classified into three categories: (1) Lambda-Shaped Attention Mechanisms, which stabilize PPL but compromise the ability to capture long-range dependencies across distant tokens Xiao et al. (2023); Han et al. (2024); Jiang et al. (2024); Li et al. (2024a); (2) Fine-Tuning on long texts, which involves training on datasets with extended positional contexts using interpolation Ding et al. (2024); Li et al. (2024b); Wu et al. (2024) or extrapolation Zhu et al. (2023); Chen et al. (2023c); Ding et al. (2024). While effective, this approach is resource-intensive and still encounters cases where positional IDs exceed its fine-tuned context window; and (3) Training-free length extrapolation methods, which include Rotary Position Embedding (RoPE) frequency interpolation techniques (e.g., Neural Tangent Kernel (NTK), Dyn-NTK, YaRN) bloc97 (2023b, a); Peng et al. (2023) and inputs rearrangement strategies (e.g., SelfExtend, ChunkLlama) Jin et al. (2024); An et al. (2024b).
However, these training-free methods exhibit significant shortcomings: (a) They rely on a global scaling factor, leading to sensitivity and inconsistent performance across both long-context and short-context tasks. (b) methods like NTK, Dyn-NTK, and YaRN suffer from attention logit outliers due to their positional embedding interpolations; and (c) SelfExtend and ChunkLlama inherently disrupt local positional relationships, compromising model performance.
To overcome these limitations, we propose Greedy Attention Logit Interpolation (GALI), a novel training-free length extrapolation method.
The innovations of GALI are twofold: 1) Greedy and Localised Interpolation: Instead of applying global scaling across all positions, GALI retains the pretrained positional IDs within the training context window, ensuring that performance on short-context inputs remains uncompromised. For tokens beyond the training context window, GALI performs interpolation at a fine-grained, token- or chunk-specific level. This greedy, localised approach eliminates redundant extrapolation, enabling stable handling of long-context tasks. 2) Logit-Level Interpolation with Positional Noise: Unlike prior work that manipulates positional embeddings, GALI operates on attention logit. It interpolates logit between valid positional pairs and injects Gaussian noise scaled to their positional interval. This design captures the oscillatory characteristics of RoPE while preventing numerical instability, resulting in robust length extrapolation.
With this, GALI explicitly addresses the shortcomings of existing methods by: (a) providing stable and superior performance on long-context tasks without compromising short-context tasks performance, eliminating the need for input-length-specific tuning, and preserving local positional information; and (b) avoiding attention logit outliers through attention logit interpolation rather than positional embedding interpolation. We conducted extensive experiments across diverse long-context benchmarks and tasks, including LongBenchBai et al. (2024), L-EvalAn et al. (2024a), and PG19Rae et al. (2019), demonstrating that GALI consistently outperforms existing state-of-the-art training-free methods. Furthermore, our analysis reveals a key insight: constraining interpolation to narrower positional intervals leads to improved performance, even on short-context tasks.
Main contributions are summarized as follows:
  • We propose Greedy Attention Logit Interpolation (GALI), a training-free method that achieves superior and stable performance across both short- and long-context tasks without any input-length-specific tuning. GALI integrates two key components: a greedy and localized position ID interpolation strategy, and a logit-level interpolation mechanism with Gaussian noise to simulate RoPE’s oscillatory behavior. These designs eliminate redundant extrapolation and attention logit outliers, enabling robust length extrapolation.
  • Our extensive evaluation on LongBench, L-Eval, and PG19 shows that GALI consistently outperforms existing training-free extrapolation methods. Further analysis reveals a key insight: constraining extrapolation to narrower positional intervals improves performance even on short-context tasks, emphasizing the importance of precise positional alignment in effective length extrapolation.
2 RelatedWork
@@
Rotary Position Embedding (RoPE): RoPE Su et al. (2024) is a technique that encodes positional information by applying rotary transformations to token embeddings, enabling relative position modeling in transformers. Given two token embeddings $\bm{x}_{m},\bm{x}_{n}\in\mathbb{R}^{l}$ as query and key corresponding to position $m$ and $n$, the projection matrix $\bm{W}_{Q},\bm{W}_{K}\in\mathbb{R}^{d\times l}$, RoPE applies a rotation to the projected token embeddings, i.e., $\bm{q}_{m}^{r}=(\bm{W}_{Q}\bm{x}_{m})e^{im\bm{\theta}}$, $\bm{k}_{n}^{r}=(\bm{W}_{K}\bm{x}_{n})e^{in\bm{\theta}}$, where $\bm{\theta}=[b^{0},b^{-2/d},\dots,b^{-2(j-1)/d}]$, $j\in[1,2,\dots,d/2]$ and $b$ is originally set to 10000. After that, the inner product between the query $\bm{q}_{m}^{r}$ and key $\bm{k}_{n}^{r}$ can be represented by the real part of ${\bm{q}_{m}^{r}}^{*}\bm{k}_{n}^{r}$, i.e.:
@@
$a(\cdot)$ is the function mapping token embeddings $\bm{x}_{m},\bm{x}_{n}$ to the attention logit, which depends on their positional interval and is irrelevant to their absolute positions. Additionally, RoPE exhibits a long-term decay as positional interval increases Su et al. (2024), as illustrated in Figure 5. Our proposed method, GALI, leverages two key properties of RoPE to achieve position interpolation and length extrapolation effectively.
Positional Out-Of-Distribution (O.O.D.): In Transformer architectures, the self-attention mechanism is inherently position-agnostic, necessitating the use of position embeddings to encode positional information for processing ordered inputs Dufter et al. (2021); Kazemnejad et al. (2024). Even in large language models (LLMs) with causal attention, explicit positional encoding through position embeddings remains the standard approach. [^1] Recent studies suggest causal attention implicitly encodes positional information, enabling performance without explicit position embeddings. However, this is beyond the scope of this paper. During inference, when LLMs encounter input sequences exceeding the maximum length seen during training, the use of unseen position IDs causes a positional out-of-distribution (O.O.D.) issue, leading to degraded performance Chen et al. (2023b); Jin et al. (2024); Xu et al. (2024). In the Rotary Position Embedding (RoPE) mechanism, extrapolating position IDs beyond the training range introduces untrained positional intervals, disrupting the attention score distribution. In contrast, position interpolation has yielded more stable attention distributions, requiring fewer fine-tuning steps. This observation has inspired subsequent interpolation-based methods Chen et al. (2023a); Xiong et al. (2023); Li et al. (2023); Ding et al. (2024); Li et al. (2024b); Wu et al. (2024), as well as training-free approaches that map position interpolation into alternative frequency dimensions in embeddings bloc97 (2023b, a); Peng et al. (2023). Recent work has explored other training-free length extrapolation techniques, such as group position IDs Jin et al. (2024) or chunk attention An et al. (2024b).
@@ 3 Method
Figure 1: The overall procedure of the proposed GALI framework. The green dashed line illustrates position ID interpolation, while the blue dashed line shows attention logit interpolation. In this example, the training context window $L_{tr}$ is 4, the chunk size $s$ is 2, the local window $L_{w}$ is 2, and the prefill length is 6. Chunks are denoted as $c_{1}$ (first chunk) and $c_{2}$ (second chunk), while $t_{7}$ represents the first generated token. The positional interval matrix $R$ incorporates $\lceil R\rceil$, $\lfloor R\rfloor$, and $R^{\prime}$, representing ceiling, floor, and modulo operations, respectively. Red numbers in the positional interval matrix indicate interpolated positional intervals and $\mathcal{N}$ represents the Gaussian noise.
We introduces Greedy Attention Logit Interpolation (GALI), a novel training-free length extrapolation method that achieves superior and consistent performance across both short- and long-context tasks without requiring any input-length-specific tuning. GALI accomplishes this through two key mechanisms: (1) a greedy and localized interpolation strategy that preserves pretrained position IDs within the training context window and only interpolates beyond it when necessary, and (2) logit-level interpolation that avoids attention logit outliers by approximating attention logit between valid relative positions. To further stabilize attention behavior, GALI adds distance-scaled Gaussian noise that simulates the oscillatory nature of RoPE. The overall process is shown in Figure 1.
@@ 3.1 Position ID Interpolation
The proposed GALI introduces a greedy and localized interpolation strategy that minimizes deviation from pretrained positional distributions. Instead of applying global scaling across all positions, as done in NTK, Dyn-NTK, YaRN, or SelfExtend, GALI retains original position IDs within the training context window and interpolates only when necessary. This fine-grained, chunk / token-wise approach avoids redundant extrapolation and eliminates the sensitivity to input length observed in prior methods.
This strategy builds on insights from Dyn-NTK bloc97 (2023a), which adjusts scaling factors based on the input length. However, Dyn-NTK still applies a global scaling factor to the entire sequence, overlooking that different tokens require different amounts of interpolation. For example, a token just beyond the training context window only needs one new position ID, while later tokens require more. Ideally, each token would have its own customized interpolation, but this is computationally expensive. GALI addresses this by grouping tokens into chunks and applying chunk-specific interpolation, avoiding global scaling.
Concretely, GALI segments the portion of the input beyond the training context window into fixed-size chunks when computing the positional interval matrix. Within each chunk, a local window of length $L_{w}$ preserves the original pretrained positional IDs. Only the remaining tokens are assigned interpolated IDs, determined by how many positions exceed the training context window, and computed to minimize disruption.
@@
In the prefill stage, given an input sequence $S=(w_{1},w_{2},\dots,w_{L_{tr}},\dots,w_{L})$ where $L_{tr}$ is the training context window size and $L$ is the input length in the prefill stage. We first divide it into chunks $C=(c_{1},c_{2},\dots,c_{L_{c}})$, where the size of $c_{1}$ is $L_{tr}$ and other chunks have a size $s$, so the $L_{c}=\lceil{\frac{L-L_{tr}}{s}}\rceil+1$. After that, we assign position IDs $S=(s_{1},s_{2},\dots,s_{L_{c}})$ for each chunk, where $s_{1}=[0,1,\dots,L_{tr}-1]$ and others are interpolated position IDs according to the following formula:
$\begin{cases}&j>1;\\ &{g}_{j}=\lceil{\frac{\sum_{i=1}^{j}{len(c_{i})}-L_{w}}{L_{tr}-L_{w}}}\rceil;\\ &{v}_{j}=1/{g}_{j};\\ &s_{j}=[0,1*{v}_{j},2*{v}_{j},\dots,({g}_{j}-1)*{v}_{j},\\ &1,\dots,L_{tr}-L_{w}-1,L_{tr}-L_{w},\dots,L_{tr}-1]\end{cases}$ (2)
Note that each chunk uses the complete pretrained positional intervals, making use of all the pretrained positional information greedily. During decoding, where tokens are generated sequentially, GALI applies the same greedy principle: each new token is treated as a single-token chunk, and its attended position IDs are interpolated based on how many attended tokens exceed the training context window.
Overall, this strategy ensures that each chunk or token reuses the full range of pretrained positional intervals when constructing the positional interval matrix. It avoids unnecessary positional distortion and removes the need for input-length-specific tuning required by global scaling methods. If the input length is within $L_{tr}$, no interpolation is applied; otherwise, the model preserves fidelity within the trained range and applies minimal-impact interpolation only to the extended portion.